# GRAM: Spatial general-purpose audio representation models for real-world applications

### v3

@@ 3.5 Ablations
To establish crucial factors for successful spatial general-purpose audio representations learning, we carried out a series of ablation experiments. For GRAM-Binaural, we trained also an encoder with a state-of-the-art 8-layer Mamba architecture (Gu & Dao, 2023; Yadav & Tan, 2024) to assess the optimal architecture choice for spatial general-purpose audio representation learning. To ensure that computational overhead and model capacity were comparable between the Transformer and Mamba encoder, we used similar parameter counts. We also tested the impact of mask type for GRAM-Binaural, comparing patch-based masking as described above to time-based masking. For time-based masking, patches were defined as $P_{i}\in\mathbb{R}^{2\times 2\times 128}$ such that they spanned the entire frequency range. For time-based masking, we used window sizes and [2, 5, 10, 25, 50, 0, 0, 0] to implement local-global attention in the decoder. For both GRAM-Binaural and GRAM-Ambisonics, we assessed the optimal ratio ($\lambda$) between simulated real-world sound scenes and clean, dry sound clips in pretraining data for $\lambda=0.0,0.25,0.5,0.75,1.0$. Finally, we examined various masking ratios [0.4, 0.6, 0.8, 0.9] and in-batch sampling factors [4, 8, 16] for GRAM-Binaural. For all ablations, GRAM-Binaural and - if applicable - GRAM-Ambisonics were trained with the exact same parameters specified above, except the masking ratio ablation, where we reduced the effective batch size from 1536 to 384.
@@ 4 Results
Generalization to localization in recorded real-world sound scenes: Table 2 shows that GRAM-Ambisonics generalizes successully to the real-world sound scenes in the TUT Sound Events 2018 dataset (Adavanne et al., 2019), obtaining a lower DoA error than other self-supervised models. GRAM-Ambisonics even outperforms supervised models such as SeldNET (Adavanne et al., 2018) (note that PILOT (Schymura et al., 2021) is sound localization model trained directly on TUT Sound Events 2018).
@@
SeldNET (Adavanne et al., 2018) 26.6
PILOT (Schymura et al., 2021) 4.2
@@
Spatial Librispeech Sarabia et al. (2023) 12.4
ELSA (Devnani et al., 2024) 15.0
@@
Figure 4: Ablation Studies. Effect of hyperparameters on HEAR and NatHEAR Performance. From left to right; (1) GRAM-Binaural downstream performance as a function of the ratio $\lambda$ between clean and naturalistic scenes in the pre-training data. (2) GRAM-Ambisonics downstream performance as a function of the ratio $\lambda$ between clean and naturalistic scenes in the pre-training data. (3) effect of masking strategy to downstream performance for GRAM-Binaural (4) comparison of Mamba and Transformer architectures on binaural training data. Important to note that architectures depicted in (4) was trained on reduced batch size (96 $\rightarrow$ 32).
@@ 4.1 Ablation Studies
Ratio of clean and naturalistic data in pretraining: Prior work on learning spatially aware audio representations from spectrograms demonstrated that pretraining on a mixture of clean and naturalistic sound scenes rather than on naturalistic sound scenes only, benefits the quality of learned representation (Devnani et al., 2024). We therefore investigated to what extent pretraining on a mixture of clean and naturalistic sound scenes affected the performance of GRAM-Binaural and GRAM-Ambisonics on HEAR and NatHEAR. The panels on the left in Figure 4 show that the performance of GRAM-Binaural on Nat-HEAR increases with lower $\lambda$, while performance on HEAR is optimal with a mixture of clean and naturalistic scenes ($\lambda$ = 0.5). GRAM-Ambisonics performed best on Nat-HEAR using a mixture of clean and naturalistic scenes ($\lambda$ = 0.5) in line with Devnani et al. (2024). In contrast, GRAM-Ambisonics performed better on HEAR with more clean data during pretraining (high $\lambda$).
@@
Encoder architecture: As shown in Figure 4, an encoder with a Transformer backbone consistently performed better than an encoder with a Mamba backbone both on clean downstream tasks (HEAR) and on naturalistic downstream tasks (NatHEAR).
@@ 5 Discussion and conclusion
Limitations and future work: Although GRAM performs well on HEAR speech tasks in comparison to other self-supervised models trained on AudioSet, we plan to train GRAM on a mixture of speech and general audio data (e.g., AudioSet) in order to assess GRAMs capability for speech learning in more detail. Furthermore, GRAM opens a path towards multi-modal spatial learning and can serve as a basis for downstream applications such as audio-visual scene representation learning (Mahmud & Marculescu, 2023), robotics Ledder et al. (2025), and audio-language representation learning (Zheng et al., 2024; Chu et al., 2023).
@@ Acknowledgments
This project received funding from the NWO Talent Program (VI.Veni.202.184; KH). We thank SURF (www.surf.nl) for the support in using the National Supercomputer Snellius.
@@
We generate our BRIRs using the simulator provided by SoundSpaces 2.0 (Chen et al., 2022a). Our hyperparameters for the simulator is depicted in Table 3.
@@ Appendix B Intensity Vectors
Following Devnani et al. (2024); Wang et al. (2022b), we extracted intensity vectors utilizing the equation below:
@@ Appendix C HEAR and Nat-HEAR Tasks
DCASE DCASE-2016 Task 2 (Mesaros et al., 2018) Event detection of overlapping office sounds in synthetic mixtures Scene Analysis
FS50K FSD50k (Fonseca et al., 2022) Multilabel, large scale audio tagging Scene Analysis
LC LibriCount (Stöter et al., 2018) Speaker Count Identification, Simulated Cocktail Party Scene Analysis
ESC-50 ESC-50 (Piczak, ) Environmental Sound Classification Environmental Sound Classification
CD Crema-D (Cao et al., 2014) Emotion Recognition Speech Analysis
VL VoxLingua107 Top10 (Valk & Alumäe, 2021) Spoken language identification Speech Analysis
SC-5 Speech Command 5h (Warden, 2018) Keyword Spotting, reduced training subset Speech Analysis
NS NSynth Pitch 5h (Engel et al., 2017) Pitch Classification, reduced training subset Pitch Classification
BO Beijing Opera (Tian et al., 2014) Classifying percussion instruments Percussion
Mri-S Mridangam Stroke (Anantapadmanabhan et al., 2013) Stroke classification in pitched percussion instruments Percussion
Mri-T Mridangam Tonic (Anantapadmanabhan et al., 2013) Tonic classification in pitched percussion instruments Percussion
@@ Appendix D Extracting GRAM embeddings for downstream tasks
We extracted GRAM embeddings for downstream evaluations by encoding embeddings for all patches $P_{1},\ldots,P_{n}$ using the GRAM encoder. We used the exact patch aggregation process as in (Niizumi et al., 2022). Audio clips were split into non-overlapping 2-second chunks and the embedded patches concatenated over time. Finally, we took the mean over the time axis to generate scene embeddings independent of the input audio duration. Finally, to evaluate GRAMs on the localization tasks, we used [CLS] embeddings of the 2-second samples, and averaged them to create scene embeddings for localization tasks.
@@ Appendix E Downstream performance metric
Similar to the procedure in SUPERB (Yang et al., 2021), let $s_{t}$ be the metric for task $t$. We then calculate the generalizability metric $\text{HEAR}_{s}(m)$, $\text{ARCH}_{s}(m)$ and $\text{NatHEAR}_{s}(m)$ for model $m$ as:
@@ Appendix F Additional Ablation Studies
Firstly, we further investigated the masking ratio, and in batch sampling as a function of HEAR and NatHEAR performance. Secondly, we investigated the localization performance in terms of mixture of naturalistic and clean audio $\lambda$.
Figure 5: Additional ablation studies. Effect of hyperparameters on HEAR and NatHEAR Performance. From left to right; (1) GRAM-Binaural downstream performance as a function of the number of in batch samples. (2) The effect of masking ratio for GRAM-Binaural. Important to note that GRAM-Binaural depicted in (2) was trained on reduced number of samples (16 $\rightarrow$ 4).
In-batch sampling: Figure 5 1 depicts that in-batch sampling helped immensely with the downstream performance on both HEAR and NatHEAR downstream. Increasing the number of in-batch samples leads to higher batch sizes with minimal computational constraints.
Masking ratio: Figure 5 2 depicts that optimal masking ratio is 0.6 for HEAR and NatHEAR performance, and higher masking ratios, such as 0.9 harms the performance.
@@ Appendix H Evaluating training efficiency
Audio-MAE (Huang et al., 2022) 512 32 3829 N/A 10s
BEATs (Chen et al., 2023) 5600 N/A N/A 1.2M 10s
SSAM (Yadav & Tan, 2024) 1024 100 2003 N/A 2s

### v4

@@ 3.5 Ablations
To establish crucial factors for successful spatial general-purpose audio representations learning, we carried out a series of ablation experiments. For GRAM-Binaural, we trained also an encoder with a state-of-the-art 8-layer Mamba architecture (mamba; audiomamba) to assess the optimal architecture choice for spatial general-purpose audio representation learning. To ensure that computational overhead and model capacity were comparable between the Transformer and Mamba encoder, we used similar parameter counts. We also tested the impact of mask type for GRAM-Binaural, comparing patch-based masking as described above to time-based masking. For time-based masking, patches were defined as $P_{i}\in\mathbb{R}^{2\times 2\times 128}$ such that they spanned the entire frequency range. For time-based masking, we used window sizes and [2, 5, 10, 25, 50, 0, 0, 0] to implement local-global attention in the decoder. For both GRAM-Binaural and GRAM-Ambisonics, we assessed the optimal ratio ($\lambda$) between simulated real-world sound scenes and clean, dry sound clips in pretraining data for $\lambda=0.0,0.25,0.5,0.75,1.0$. Finally, we examined various masking ratios [0.4, 0.6, 0.8, 0.9] and in-batch sampling factors [4, 8, 16] for GRAM-Binaural. For all ablations, GRAM-Binaural and - if applicable - GRAM-Ambisonics were trained with the exact same parameters specified above, except the masking ratio ablation, where we reduced the effective batch size from 1536 to 384.
@@ 4 Results
Generalization to localization in recorded real-world sound scenes: Table 2 shows that GRAM-Ambisonics generalizes successully to the real-world sound scenes in the TUT Sound Events 2018 dataset (tut), obtaining a lower DoA error than other self-supervised models. GRAM-Ambisonics even outperforms supervised models such as SeldNET (adavanne2018sound) (note that PILOT (pilot) is sound localization model trained directly on TUT Sound Events 2018).
@@
SeldNET (adavanne2018sound) 26.6
PILOT (pilot) 4.2
@@
Spatial Librispeech spatial-libri 12.4
ELSA (elsa) 15.0
@@
Figure 4: Ablation Studies. Effect of hyperparameters on HEAR and Nat-HEAR Performance. From left to right; (1) GRAM-Binaural downstream performance as a function of the ratio $\lambda$ between clean and naturalistic scenes in the pre-training data. (2) GRAM-Ambisonics downstream performance as a function of the ratio $\lambda$ between clean and naturalistic scenes in the pre-training data. (3) effect of masking strategy to downstream performance for GRAM-Binaural (4) comparison of Mamba and Transformer architectures on binaural training data. Important to note that architectures depicted in (4) was trained on reduced batch size (96 $\rightarrow$ 32).
@@ 4.1 Ablation Studies
Ratio of clean and naturalistic data in pretraining: Prior work on learning spatially aware audio representations from spectrograms demonstrated that pretraining on a mixture of clean and naturalistic sound scenes rather than on naturalistic sound scenes only, benefits the quality of learned representation (elsa). We therefore investigated to what extent pretraining on a mixture of clean and naturalistic sound scenes affected the performance of GRAM-Binaural and GRAM-Ambisonics on HEAR and Nat-HEAR. The panels on the left in Figure 4 show that the performance of GRAM-Binaural on Nat-HEAR increases with lower $\lambda$, while performance on HEAR is optimal with a mixture of clean and naturalistic scenes ($\lambda$ = 0.5). GRAM-Ambisonics performed best on Nat-HEAR using a mixture of clean and naturalistic scenes ($\lambda$ = 0.5) in line with elsa. In contrast, GRAM-Ambisonics performed better on HEAR with more clean data during pretraining (high $\lambda$).
@@
Encoder architecture: As shown in Figure 4, an encoder with a Transformer backbone consistently performed better than an encoder with a Mamba backbone both on clean downstream tasks (HEAR) and on naturalistic downstream tasks (Nat-HEAR).
@@ 5 Discussion and conclusion
Limitations and future work: Although GRAM performs well on HEAR speech tasks in comparison to other self-supervised models trained on AudioSet, we plan to train GRAM on a mixture of speech and general audio data (e.g., AudioSet) in order to assess GRAMs capability for speech learning in more detail. Furthermore, GRAM opens a path towards multi-modal spatial learning and can serve as a basis for downstream applications such as audio-visual scene representation learning (clip), robotics ledder2025audio, and audio-language representation learning (spatialast; chu2023qwen).
@@ Acknowledgments
This project received funding from the NWO Talent Program (VI.Veni.202.184; KH). This work used the Dutch national e-infrastructure with the support of the SURF Cooperative using grant no. EINF-12218. We would like to thank Robert Jan Schlimbach from the Snellius team for helpful discussions and their help with high performance cluster utilization.
@@
We generate our BRIRs using the simulator provided by SoundSpaces 2.0 (chen22soundspaces2). Our hyperparameters for the simulator is depicted in Table 3.
@@ Appendix B Intensity Vectors
Following elsa; ambisonic_IVs, we extracted intensity vectors utilizing the equation below:
@@ Appendix C HEAR and Nat-HEAR Tasks
DCASE DCASE-2016 Task 2 (Mesaros2018_TASLP) Event detection of overlapping office sounds in synthetic mixtures Scene Analysis
FS50K FSD50k (fsd50k) Multilabel, large scale audio tagging Scene Analysis
LC LibriCount (libricount) Speaker Count Identification, Simulated Cocktail Party Scene Analysis
ESC-50 ESC-50 (esc50) Environmental Sound Classification Environmental Sound Classification
CD Crema-D (cremad) Emotion Recognition Speech Analysis
VL VoxLingua107 Top10 (voxling) Spoken language identification Speech Analysis
SC-5 Speech Command 5h (warden2018speechcommandsdatasetlimitedvocabulary) Keyword Spotting, reduced training subset Speech Analysis
NS NSynth Pitch 5h (nsynth2017) Pitch Classification, reduced training subset Pitch Classification
BO Beijing Opera (Beijingop) Classifying percussion instruments Percussion
Mri-S Mridangam Stroke (mrist) Stroke classification in pitched percussion instruments Percussion
Mri-T Mridangam Tonic (mrist) Tonic classification in pitched percussion instruments Percussion
@@ Appendix D Extracting GRAM embeddings for downstream tasks
We extracted GRAM embeddings for downstream evaluations by encoding embeddings for all patches $P_{1},\ldots,P_{n}$ using the GRAM encoder. We used the exact patch aggregation process as in (masked-modelling). Audio clips were split into non-overlapping 2-second chunks and the embedded patches concatenated over time. Finally, we took the mean over the time axis to generate scene embeddings independent of the input audio duration. Finally, to evaluate GRAMs on the localization tasks, we used [CLS] embeddings of the 2-second samples, and averaged them to create scene embeddings for localization tasks.
@@ Appendix E Downstream performance metric
Similar to the procedure in SUPERB (superb), let $s_{t}$ be the metric for task $t$. We then calculate the generalizability metric $\text{HEAR}_{s}(m)$, $\text{ARCH}_{s}(m)$ and $\text{Nat-HEAR}_{s}(m)$ for model $m$ as:
@@ Appendix F Additional Ablation Studies
Firstly, we further investigated the masking ratio, and in batch sampling as a function of HEAR and Nat-HEAR performance. Secondly, we investigated the localization performance in terms of mixture of naturalistic and clean audio $\lambda$.
Figure 5: Additional ablation studies. Effect of hyperparameters on HEAR and Nat-HEAR Performance. From left to right; (1) GRAM-Binaural downstream performance as a function of the number of in batch samples. (2) The effect of masking ratio for GRAM-Binaural. Important to note that GRAM-Binaural depicted in (2) was trained on reduced number of samples (16 $\rightarrow$ 4).
In-batch sampling: Figure 5 1 depicts that in-batch sampling helped immensely with the downstream performance on both HEAR and Nat-HEAR downstream. Increasing the number of in-batch samples leads to higher batch sizes with minimal computational constraints.
Masking ratio: Figure 5 2 depicts that optimal masking ratio is 0.6 for HEAR and Nat-HEAR performance, and higher masking ratios, such as 0.9 harms the performance.
@@ Appendix H Evaluating training efficiency
Audio-MAE (audiomae) 512 32 3829 N/A 10s
BEATs (beats) 5600 N/A N/A 1.2M 10s
SSAM (audiomamba) 1024 100 2003 N/A 2s